7. Sulfuryl chloride, SO2Cl2, is a highly reactive gaseous compound. When heated, it decomposes as follows: SO2Cl2 (g) ⇌ SO2 (g) + Cl2 (g). This decomposition is endothermic. A sample of 3.509 grams of SO2Cl2 is placed in an evacuated 1.00-L bulb and the temperature is raised to 375 K.
a. What would be the pressure (in atm) in the bulb if no dissociation of the SO2Cl2(g) occurred?
b. When the system has come to equilibrium at 375 K, the total pressure in the bulb is found to be 1.43 atm. Calculate the partial pressures of SO2, Cl2, and SO2Cl2 at equilibrium.
c. Give the expression for the equilibrium constant (either Kp or KC) for the decomposition of
SO2Cl2(g) at 375 K. Calculate the value of the equilibrium constant you have selected.
d. If the temperature were raised to 500 K, what effect would this have on the equilibrium
constant? Explain in 1-2 sentences.
1
Expert's answer
2020-05-11T14:23:27-0400

pV=nRT. n=m/Mr = 3.509 / 135 = 0.026 moles.

a) p=nRT/V=0.026 x 0.082 x 375/1 = 0.8 atm.

b) SO2Cl2 (g) ⇌ SO2 (g) + Cl2 (g).

1.43 - 0.8 = 0.63

2-1 =1.

0.63 x 0.026 = 0,01638. 0.026 -0,01638= 0.00962 moles of SO2CL2.

0.01638 moles of SO2 and 0.01638 moles of Cl2.

0.01638 + 0.01638 + 0.00962 = 0.04238 moles total.

1.43 x 0.01638 / 0.04238 = 0.5526 atm of SO2 and Cl2.

1.43 x 0.00962 / 0.04238 = 0.3246 atm of SO2CL2.

c) 0.01638 x 0.01638/ 0.00962=0,0278902703

d) will react turn to right.